OpenStudy (anonymous):

What is the solution to the equation log4 x + log4 8 = 3 ? x = one–half x = 2 x = 8 x = 32

OpenStudy (acacia):

log4 x + log4 8 = 3 log4 8x = 3 8x = 4^3 8x = 64 x = 8

OpenStudy (nurali):

log4 x + log4 8 = 3 log4 8x = 3 .. as log a + log b = log ab 8x = 4 ^3 raise to the power of 4 on both side and (a^loga x = x) x = 64/8 = 8